OpenStudy (anonymous):

a.The Bread Alone Bakery has a daily overhead of $90. It costs $0.60 to bake each loaf of bread, and the bread sells for $1.50 per loaf. a. Write an equation for the cost, C, in terms of the number of loaves, x. C= b. Write an equation for the revenue, R, in terms of the number of loaves, x. R=

OpenStudy (anonymous):

C=.9x 90=.9x /.9 /.9 x=150 R=1.5x - .9x R=1.5(150) - .9(150) R=225 - 90 R= 135 ...but in my economics class, revenue is simply quantity sold times price sold for. In that case, total revenue would be $225

OpenStudy (anonymous):

Whoops, a little error in math. I flipped the .6 to .9

OpenStudy (anonymous):

C=.6x 90=.6x /.6 /.6 x=150 R=1.5x - .6x R=1.5(150) - .6(150) R=225 - 90 R= 135 The math was actually right, just flipped the number when typing it up. Revenue will be $135 or $225 depending on definition of Revenue. Makes sense?

OpenStudy (anonymous):

YES11 THANK YOU:)

OpenStudy (anonymous):

Just a thought, the overhead cost of $90 could be the rent or something or another... and not the cost of daily bread baking. The question is a unclear on that. Hope that helps!